solve by subsitution method can you assit me please
0.05x-0.01y= -0.1
0.28x+0.13y=10.6

Answer by Alan3354(69443)

solve by subsitution method can you assit me please
0.05x-0.01y= -0.1 x13 --> 0.65x - 0.13y = -1.3 --> 0.13y = 0.65x + 1.3
0.28x+0.13y=10.6
-------------------
Sub for 0.13y in the 2nd eqn
--------
0.28x + 0.65x + 1.3 = 10.6
0.93x = 9.3
x = 10
----------
0.05*10 - 0.01y = -0.1
-0.01y = -0.6
y = 60